OpenStudy (anonymous):

Identify the transformations required to transform y=x^2-3x-10 into y=x^2

OpenStudy (anonymous):

fak fak fak fak fak fak

OpenStudy (anonymous):

Hi Welcome to Openstudy :) sorry for the first response you get First thing you have to do is to transform the equation \(y=x^2-3x-10\) into vertex form \(y=a(x+h)^2+k\) can you do that? then after that, refer to this image that i attached and you can be able to determine the transformations of the function